a SHIPMENT OF 18 CARS some weighing 3000 pounds and others 5000 each has a toal weight of 30 tones find the number of each kind of car
Answer by checkley77(12844)

X+Y=18 OR X=18-Y
3,000X+5,000Y=30*2,000
3,000(18-Y)+5,000Y=60,000
54,000-3,000Y+5,000Y=60,000
2,000Y=60,000-54,000
2,000Y=6,000
Y=6,000/2,000
Y=3 5,000 POUND CARS ARE SHIPPED.
18-3=15 3,000 POUND CARS ARE SHIPPED.
PROOF:
3,000*15+5,000*3=60,000
45,000-15,000=60,000
60,000=60,000